# Are Multimodal Embeddings Truly Beneficial for Recommendation? A Deep Dive into Whole vs. Individual Modalities

### v1

@@ Abstract.
Multimodal recommendation (MMRec) has emerged as a mainstream paradigm, typically leveraging text and visual embeddings extracted from pre-trained models such as Sentence-BERT, Vision Transformers, and ResNet. This approach is founded on the intuitive assumption that incorporating multimodal embeddings can enhance recommendation performance. However, despite its popularity, this assumption lacks comprehensive empirical verification. This presents a critical research gap. To address it, we pose the central research question of this paper: Are multimodal embeddings truly beneficial for recommendation?
To answer this question, we conduct a large-scale empirical study examining the role of text and visual embeddings in modern MMRec models, both as a whole and individually. Specifically, we pose two key research questions: (1) Do multimodal embeddings as a whole improve recommendation performance? (2) Is each individual modality—text and image—useful when used alone? To isolate the effect of individual modalities—text or visual—we employ a modality knockout strategy by setting the corresponding embeddings to either constant values or random noise. To ensure the scale and comprehensiveness of our study, we evaluate 14 widely used state-of-the-art MMRec models. Our findings reveal that: (1) multimodal embeddings generally enhance recommendation performance—particularly when integrated through more sophisticated graph-based fusion models. Surprisingly, commonly adopted baseline models with simple fusion schemes, such as VBPR and BM3, show only limited gains. (2) The text modality alone achieves performance comparable to the full multimodal setting in most cases, whereas the image modality alone does not. These results offer foundational insights and practical guidance for the MMRec community. We will release our code and datasets to facilitate future research.
@@ 1. Introduction
Multimodal recommendation (MMRec) has emerged as a powerful paradigm in modern recommender systems, enabling the integration of diverse modalities such as text, image, video, and audio. By leveraging pretrained encoders such as Sentence-BERT, Vision Transformers (ViT), and ResNet, recent state-of-the-art models—including BM3 (Zhou et al., 2023a), FREEDOM (Zhou and Shen, 2023), and SMORE (Ong and Khong, 2025)—have demonstrated impressive performance across various recommendation tasks. These models are built on the intuitive yet widely accepted assumption that incorporating multimodal embeddings provides richer user-item representations and thus improves recommendation accuracy.
@@
Although extensive research has been conducted in the MMRec domain, evaluating the true effectiveness of multimodal embeddings remains challenging due to several factors (Beel et al., 2016): (1) most studies emphasize novel architectural designs or training strategies rather than directly examining the intrinsic quality of the embeddings; (2) evaluation protocols vary widely—including different datasets, metrics, and data-splitting strategies—making cross-study comparisons largely inconclusive; (3) reported results often reflect a combination of influences such as model architecture, optimization techniques, and training heuristics, which obscures the isolated contribution of the embeddings themselves.
To overcome the aforementioned limitations and answer the central question, we conduct a large-scale empirical study within a controlled and consistent experimental environment, focusing specifically on the role of multimodal embeddings in recommendation. We adopt a modality knockout strategy, where the embedding of a given modality is replaced with either constant vectors (e.g., all zeros), or random noise. This simple yet effective approach allows us to isolate the contribution of each modality without introducing confounding architectural factors. Our investigation centers on two sub-questions: (1) Do multimodal embeddings as a whole enhance recommendation performance? and (2) Is each individual modality—text or image—beneficial when used in isolation? To address these questions, we apply our methodology to 14 widely-used MMRec models, spanning a diverse set of fusion techniques and architectural families, ensuring a comprehensive and representative evaluation.
Our findings yield several notable insights. First, we confirm that leveraging multimodal embeddings generally enhances recommendation performance in most evaluated models equipped with sophisticated fusion mechanisms. However, we also observe some surprising results in methods employing simpler fusion strategies, such as VBPR and BM3, which are commonly used as baselines in MMRec. Second, we find that the text modality alone often performs comparably to the full multimodal setting, highlighting its dominant role. In contrast, the image modality alone provides only limited performance gains, and in many cases, its removal does not significantly degrade accuracy. The visualization of both text and image embeddings shows that text embeddings are more compact, whereas image embeddings are more dispersed, which may make them harder to learn.
In addition to these findings, we also summarize a set of practical techniques that help boost the performance of classic MMRec baselines, especially in resource-constrained or modality-limited scenarios.
@@
  • We present the first large-scale and systematic analysis of the role of multimodal embeddings in MMRec, evaluating 14 state-of-the-art models.
@@
  • We provide practical insights and a set of additional empirical findings that serve as actionable guidance for future MMRec model development and evaluation.
@@ 2. Related Work
Reproducibility Study of Recommender Systems. In the field of recommender systems, differences in datasets, evaluation protocols, and experimental settings make reproducibility studies particularly important (Beel et al., 2016). Rather than proposing new methodologies, these studies re-examine widely adopted baselines to draw fundamental insights for future research. For example, Ferrari Dacrema et al. (2019) reveals that many state-of-the-art neural network-based recommendation methods perform even worse than simple nearest-neighbor techniques, Rendle et al. (2020) revisits matrix factorization in comparison with neural collaborative filtering, and Petrov and Macdonald (2022) conducts a reproducibility study on the BERT4Rec model (Sun et al., 2019). Ji et al. (2023) presents a critical study on the offline evaluation of recommender systems. In the recent modality-based recommendation domain, many works (Li et al., 2023; Yuan et al., 2023; Fu et al., 2024c; Ni et al., 2023; Zhang et al., 2024a) compare advanced sequential recommendation models with heavy end-to-end encoders against simple ID embedding models, demonstrating that ID-based models still achieve strong performance. Although these studies do not propose novel methodologies, they provide valuable insights that have had a significant impact on the research community. Each serves as a cornerstone of the field, inspiring our study on recent state-of-the-art multimodal recommendation models.
@@
Multimodal recommendation (MMRec). MMRec (Liu et al., 2024; Zhou et al., 2023b; Fu et al., 2024a) leverages auxiliary information from different modalities, primarily text, and images, to enrich item representations. Early work like VBPR (He and McAuley, 2016) extended the classic Bayesian Personalized Ranking (BPR) framework by incorporating pre-extracted visual features into the matrix factorization model, demonstrating that visual style can influence user preference. With the rise of Graph Neural Networks (GNNs), which excel at capturing high-order connectivity, graph-based models became the dominant paradigm. MMGCN (Wei et al., 2019) was a pioneering work that constructed separate user-item bipartite graphs for each modality, learning modality-specific representations through graph convolution. Following this, GRCN (Wei et al., 2020) proposed refining the user-item graph by using multimodal features to identify and down-weight potential noisy (false-positive) interactions. More recent models like LATTICE (Zhang et al., 2021) and FREEDOM (Zhou and Shen, 2023) introduced an explicit item-item semantic graph built from multimodal features to capture latent item relationships, which is then combined with the user-item interaction graph. LGMRec (Guo et al., 2024) further extended this by learning both local and global user interests, using a hypergraph to model global dependencies. Other approaches, such as BM3 (Zhou et al., 2023a), adopt a self-supervised learning paradigm. Its core innovation lies in bootstrapping latent representations by generating contrastive views through a simple dropout mechanism in the embedding space, thereby eliminating the need for explicit and computationally expensive graph augmentation or construction.
Despite the growing architectural complexity and the introduction of novel designs, a common thread across these works is the implicit assumption that multimodal embeddings enhance the final recommendation performance. Limited research has empirically examined this core assumption. To the best of our knowledge, we are the first to conduct a large-scale, systematic empirical investigation of multimodal embeddings.
@@ 3. Preliminaries
Problem Definition. This paper investigates the role of multimodal embeddings in multimodal recommendation (MMRec). Let $U$ denote the set of users and $I$ the set of items. In the multimodal recommendation setting, each item $i\in I$ is associated with both visual and textual embeddings. Specifically:
@@ 4.1. Datasets
Item image features were obtained by using the pre-extracted vectors extracted by Resnet-50 (He et al., 2016) provided in the benchmark; these serve as the baseline visual embeddings. Item text features typically come from product titles or descriptions, encoded via Sentence-BERT (Reimers and Gurevych, 2019).
4.2. MMRec Models
@@ 4.3. Experimental Protocol
Figure 1. Comparing 14 MMRec Models under Modality Knockouts on Baby, Clothing, and Sports Datasets.
@@ 5.1. Multimodality as a Whole (RQ1)
(Answer to RQ1) This demonstrates that the joint use of visual and textual features as a whole provides crucial information that substantially improves recommendation performance in most MMRec models with sophisticated fusion. Interestingly, common baselines with simpler fusion strategies, such as VBPR and BM3, exhibit surprisingly stable performance, suggesting that the incorporated modality information is not beneficial in these models.
@@ 6.2. Case Study: VBPR’s Failure to Utilize Modality Features
To investigate this phenomenon more deeply, we selected the most common MMRec baseline VBPR for a more granular analysis. By systematically manipulating the numerical distribution of the replaced embeddings, we observed the following behaviors:
@@
This serves as a critical insight: when evaluating multimodal models, merely demonstrating superior performance over a baseline (e.g., VBPR ¿ BPR) is insufficient proof of efficacy. Such performance gains may originate from architectural artifacts rather than genuine multimodal learning.
@@ 7. Conclusion and Future Work
This study critically examines whether current multimodal recommendation (MMRec) models truly leverage all available modality information. Our analysis first reveals that most existing MMRec models primarily rely on multimodal embeddings, particularly those equipped with sophisticated fusion mechanisms. Surprisingly, some classic baselines with simple fusion strategies, such as VBPR and BM3, exhibit unexpectedly stable performance even when their multimodal embeddings are replaced with meaningless vectors. A closer investigation of individual modalities further shows that text plays a dominant role: most models maintain comparable performance even when image embeddings are removed. We also discuss models whose behavior deviates from this general trend.
These findings have important implications for the MMRec community. We advocate for modality ablation studies to become a standard evaluation protocol, ensuring that reported improvements over baselines are transparent and credible. Moreover, a deeper understanding of how multimodal embeddings influence MMRec performance would greatly benefit practitioners in effectively deploying these models. For future model design, researchers should move beyond simple fusion and explore integration strategies where modalities actively influence the construction of user–item and item–item graphs during intermediate computations, supported by training techniques that prevent signal overshadowing.

### v2

@@ Abstract.
Multimodal recommendation has emerged as a mainstream paradigm, typically leveraging text and visual embeddings extracted from pre-trained models such as Sentence-BERT, Vision Transformers, and ResNet. This approach is founded on the intuitive assumption that incorporating multimodal embeddings can enhance recommendation performance. However, despite its popularity, this assumption lacks comprehensive empirical verification. This presents a critical research gap. To address it, we pose the central research question of this paper: Are multimodal embeddings truly beneficial for recommendation?
To answer this question, we conduct a large-scale empirical study examining the role of text and visual embeddings in modern multimodal recommendation models, both as a whole and individually. Specifically, we pose two key research questions: (1) Do multimodal embeddings as a whole improve recommendation performance? (2) Is each individual modality—text and image—useful when used alone?
To isolate the effect of individual modalities—text or visual—we employ a modality knockout strategy by setting the corresponding embeddings to either constant values or random noise. To ensure the scale and comprehensiveness of our study, we evaluate 14 widely used state-of-the-art multimodal recommendation models. Our findings reveal that: (1) multimodal embeddings generally enhance recommendation performance—particularly when integrated through more sophisticated graph-based fusion models. Surprisingly, commonly adopted baseline models with simple fusion schemes, such as VBPR and BM3, show only limited gains. (2) The text modality alone achieves performance comparable to the full multimodal setting in most cases, whereas the image modality alone does not. These results offer foundational insights and practical guidance for the multimodal recommendation community. We will release our code and datasets to facilitate future research.
@@ 1. Introduction
Multimodal recommendation has emerged as a powerful paradigm in modern recommender systems, enabling the integration of diverse modalities such as text, image, video, and audio. By leveraging pretrained encoders such as Sentence-BERT, Vision Transformers (ViT), and ResNet, recent state-of-the-art models—including BM3 (Zhou et al., 2023a), FREEDOM (Zhou and Shen, 2023), and SMORE (Ong and Khong, 2025)—have demonstrated impressive performance across various recommendation tasks. These models are built on the intuitive yet widely accepted assumption that incorporating multimodal embeddings provides richer user-item representations and thus improves recommendation accuracy.
@@
Although extensive research has been conducted in the multimodal recommendation domain, evaluating the true effectiveness of multimodal embeddings remains challenging due to several factors (Beel et al., 2016): (1) most studies emphasize novel architectural designs or training strategies rather than directly examining the intrinsic quality of the embeddings; (2) evaluation protocols vary widely—including different datasets, metrics, and data-splitting strategies—making cross-study comparisons largely inconclusive; (3) reported results often reflect a combination of influences such as model architecture, optimization techniques, and training heuristics, which obscures the isolated contribution of the embeddings themselves.
To overcome the aforementioned limitations and answer the central question, we conduct a large-scale empirical study within a controlled and consistent experimental environment, focusing specifically on the role of multimodal embeddings in recommendation. We adopt a modality knockout strategy, where the embedding of a given modality is replaced with either constant vectors (e.g., all zeros), or random noise. This simple yet effective approach allows us to isolate the contribution of each modality without introducing confounding architectural factors. Our investigation centers on two sub-questions: (1) Do multimodal embeddings as a whole enhance recommendation performance? and (2) Is each individual modality—text or image—beneficial when used in isolation? To address these questions, we apply our methodology to 14 widely-used multimodal recommendation models, spanning a diverse set of fusion techniques and architectural families, ensuring a comprehensive and representative evaluation.
Our findings yield several notable insights. First, we confirm that leveraging multimodal embeddings generally enhances recommendation performance in most evaluated models equipped with sophisticated fusion mechanisms. However, we also observe some surprising results in methods employing simpler fusion strategies, such as VBPR and BM3, which are commonly used as baselines in multimodal recommendation. Second, we find that the text modality alone often performs comparably to the full multimodal setting, highlighting its dominant role. In contrast, the image modality alone provides only limited performance gains, and in many cases, its removal does not significantly degrade accuracy. The visualization of both text and image embeddings shows that text embeddings are more compact, whereas image embeddings are more dispersed, which may make them harder to learn.
In addition to these findings, we also summarize a set of practical techniques that help boost the performance of classic multimodal recommendation baselines, especially in resource-constrained or modality-limited scenarios.
@@
  • We present the first large-scale and systematic analysis of the role of multimodal embeddings in multimodal recommendation, evaluating 14 state-of-the-art models.
@@
  • We provide practical insights and a set of additional empirical findings that serve as actionable guidance for future multimodal recommendation model development and evaluation.
@@ 2. Related Work
Reproducibility Study of Recommender Systems In the field of recommender systems, differences in datasets, evaluation protocols, and experimental settings make reproducibility studies particularly important (Beel et al., 2016). Rather than proposing new methodologies, these studies re-examine widely adopted baselines to draw fundamental insights for future research. For example, Ferrari Dacrema et al. (2019) reveals that many state-of-the-art neural network-based recommendation methods perform even worse than simple nearest-neighbor techniques, Rendle et al. (2020) revisits matrix factorization in comparison with neural collaborative filtering, and Petrov and Macdonald (2022) conducts a reproducibility study on the BERT4Rec model (Sun et al., 2019). Ji et al. (2023) presents a critical study on the offline evaluation of recommender systems. In the recent modality-based recommendation domain, many works (Li et al., 2023; Yuan et al., 2023; Fu et al., 2024c; Ni et al., 2023; Zhang et al., 2024a) compare advanced sequential recommendation models with heavy end-to-end encoders against simple ID embedding models, demonstrating that ID-based models still achieve strong performance. Although these studies do not propose novel methodologies, they provide valuable insights that have had a significant impact on the research community. Each serves as a cornerstone of the field, inspiring our study on recent state-of-the-art multimodal recommendation models.
@@
Multimodal recommendation. Multimodal recommendation (Liu et al., 2024; Zhou et al., 2023b; Fu et al., 2024a) leverages auxiliary information from different modalities, primarily text, and images, to enrich item representations. Early work like VBPR (He and McAuley, 2016) extended the classic Bayesian Personalized Ranking (BPR (Rendle et al., 2009)) framework by incorporating pre-extracted visual features into the matrix factorization model, demonstrating that visual style can influence user preference. With the rise of Graph Neural Networks (GNNs), which excel at capturing high-order connectivity, graph-based models became the dominant paradigm. MMGCN (Wei et al., 2019) was a pioneering work that constructed separate user-item bipartite graphs for each modality, learning modality-specific representations through graph convolution. Following this, GRCN (Wei et al., 2020) proposed refining the user-item graph by using multimodal features to identify and down-weight potential noisy (false-positive) interactions. More recent models like LATTICE (Zhang et al., 2021) and FREEDOM (Zhou and Shen, 2023) introduced an explicit item-item semantic graph built from multimodal features to capture latent item relationships, which is then combined with the user-item interaction graph. LGMRec (Guo et al., 2024) further extended this by learning both local and global user interests, using a hypergraph to model global dependencies. Other approaches, such as BM3 (Zhou et al., 2023a), adopt a self-supervised learning paradigm. Its core innovation lies in bootstrapping latent representations by generating contrastive views through a simple dropout mechanism in the embedding space, thereby eliminating the need for explicit and computationally expensive graph augmentation or construction.
Concurrent work. Concurrent to ours, Pomo et al. (Pomo et al., 2025) examine whether multimodal recommendation gains stem from semantics or capacity, introducing LVLM(Large Vision–Language Model)-derived, multimodal-by-design embeddings via structured VQA prompting with EOS-based extraction. Meanwhile, Zhou et al. (Zhou et al., 2025) propose an evaluation framework and benchmark multimodal recommendation models across tasks, stages, and integration strategies to assess modality contributions. By contrast, our work diagnoses reliance via Modality Knockout: we replace the target modality with constants/noise during both training and inference while preserving the original model architecture, isolating data-signal effects and enabling consistent, model-agnostic comparison.
Despite increasing architectural complexity, many methods implicitly assume that multimodal embeddings improve recommendation performance, while rigorous tests of this assumption remain limited. To our knowledge, ours is a pioneering effort in large-scale, systematic investigations that isolate modality information via architecture-preserving knockouts, providing clear evidence of when—and to what extent—text and image embeddings matter.
@@ 3. Preliminaries
Problem Definition. This paper investigates the role of multimodal embeddings in multimodal recommendation. Let $U$ denote the set of users and $I$ the set of items. In the multimodal recommendation setting, each item $i\in I$ is associated with both visual and textual embeddings. Specifically:
@@ 4.1. Datasets
Item image features were obtained by using the pre-extracted vectors extracted by ResNet-50 (He et al., 2016) provided in the benchmark; these serve as the baseline visual embeddings. Item text features typically come from product titles or descriptions, encoded via pre-trained all-MiniLM-L6-v2 model (Reimers and Gurevych, 2020).
4.2. Models Implemented via the MMRec Toolbox
@@ 4.3. Experimental Protocol
Figure 1. Comparing 14 Multimodal Recommendation Models under Modality Knockouts on Baby, Clothing, and Sports Datasets.
@@ 5.1. Multimodality as a Whole (RQ1)
(Answer to RQ1) This demonstrates that the joint use of visual and textual features as a whole provides crucial information that substantially improves recommendation performance in most multimodal recommendation models with sophisticated fusion. Interestingly, common baselines with simpler fusion strategies, such as VBPR and BM3, exhibit surprisingly ”stable” performance, suggesting that the incorporated modality information is not beneficial in these models.
@@ 6.2. Case Study: VBPR’s Failure to Utilize Modality Features
To investigate this phenomenon more deeply, we selected the most common multimodal recommendation baseline VBPR for a more granular analysis. By systematically manipulating the numerical distribution of the replaced embeddings, we observed the following behaviors:
@@
This serves as a critical insight: when evaluating multimodal models, merely demonstrating superior performance over a baseline (e.g., VBPR >> BPR) is insufficient proof of efficacy. Such performance gains may originate from architectural artifacts rather than genuine multimodal learning.
@@ 7. Conclusion and Future Work
This study critically examines whether current multimodal recommendation models truly leverage all available modality information. Our analysis first reveals that most existing multimodal recommendation models primarily rely on multimodal embeddings, particularly those equipped with sophisticated fusion mechanisms. Surprisingly, some classic baselines with simple fusion strategies, such as VBPR and BM3, exhibit unexpectedly stable performance even when their multimodal embeddings are replaced with meaningless vectors. A closer investigation of individual modalities further shows that text plays a dominant role: most models maintain comparable performance even when image embeddings are removed. We also discuss models whose behavior deviates from this general trend.
These findings have important implications for the multimodal recommendation community. We advocate for modality ablation studies to become a standard evaluation protocol, ensuring that reported improvements over baselines are transparent and credible. Moreover, a deeper understanding of how multimodal embeddings influence multimodal recommendation performance would greatly benefit practitioners in effectively deploying these models. For future model design, researchers should move beyond simple fusion and explore integration strategies where modalities actively influence the construction of user–item and item–item graphs during intermediate computations, supported by training techniques that prevent signal overshadowing.